The following selected information was extracted from the 20x1 accounting records of Lone Oak Products:

Raw material purchases $ 175,000
Direct labor 255,000
Indirect labor 107,000
Selling and administrative salaries 133,000
Building depreciation* 80,000
Other selling and administrative expenses 195,000
Other factory costs 343,000
Sales revenue ($130 per unit) 1,495,000

*Seventy-five percent of the company's building was devoted to production activities; the remaining 25 percent was used for selling and administrative functions.


Inventory data:

January 1 December 31
Raw material $

16,000

$

18,200

Work in process

35,800

62,100
Finished goods*

111,100

97,900

*The January 1 and December 31 finished-goods inventory consisted of 1,350 units and 1,190 units, respectively.

Calculate:
A) Manufacturing Overhead:

B) Cost of Goods Manufactured:

C) Cost of Good Sold

D) Net Income

E) Number of Complete Units
Answer #1

a) Manufacturing overhead

Indirect labor 107000
Building depreciation (80000*75%) 60000
Other factory cost 343000
Manufacturing overhead 510000

b) Direct material used = 16000+175000-18200 = 172800

Total manufacturing cost = 172800+255000+510000 = 937800

Cost of goods manufactured = 35800+937800-62100 = 911500

c) Cost of goods sold = 111100+911500-97900 = 924700

d) Net income :

Sales 1495000
Cost of goods sold -924700
Gross profit 570300
Selling and administrative expense (133000+20000+195000) 348000
Net income 222300

f) No of unit completed = 1190+11500-1350 = 11340 Units